OpenStudy (smurfy14):

Two values of a linear function f(x) are f(4)=2 and f(-5)=-1. Find the equations of f(x).

OpenStudy (anonymous):

this is another way of stating that the points (4,2) and (-5,-1) are on the line. our slope is (-1-2)/(-5-4)=1/3 point slope form f(x)-2=(1/3)*(x-4) f(x)=(1/3)x-4/3+2 f(x)=(1/3)x+2/3